You have a piggy bank containing a total of 86 coins in dimes and quarters. If the piggy bank contains $15.80, how many dimes are there in the piggy bank?
Answer by addingup(3677)

d+q = 86 and therefore q = 86-d
0.10d+0.25q = 15.80 Substitute for q:
0.10d+0.25(86-d) = 15.80
0.10d+21.50-0.25d = 15.80 Add d on left and subtract 21.50 on both sides:
-0.15d = -5.70 divide both sides by -0.15 and remember -/- = +
d = 38
and
q = 86-38 = 48
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
d: 38(0.10)= 3.80
q: 48(0.25)= 12
____________ -----
Total....... 15.80 We have the right answer
J